Question 1 2 pts hat is the present value of $100 ten years from now if the simple rate of interest is0.01? Round your final answer to two decimal places. Question 2 2 pts A deposit of $17 is made at the beginning of each year for 15 years, what is the future value of all deposits if simple rate of interest is 0.01? Round your final answer to two decimal places. Question 3 2 pts In how many years will $100 become $320 if simple rate of interest is 11%?
Answer #1

(1) Let the Present Value of $ 100 be $ P

Simple Interest Rate = 0.01 or 1 %

Total Interest Generated = (P x 1 x 10) / 100 = $ 0.1 P

Total Value of Interest Plus Original Principal = P + 0.1P = $ 100

Therefore, 1.1 P = $ 100

P = 100 / 1.1 = $ 90.909 ~ $ 90.91

NOTE: Please raise separate queries for solutions to the remaining unrelated question as one query is restricted to the solution of only one complete question (with a maximum of four sub-parts)